OpenStudy (anonymous):

Ally is wearing a witch’s costume for Halloween. Her hat is in the shape of a cone with a diameter of 7 inches and a height of 18 inches. What is the volume of Ally’s hat? Use for pi.

OpenStudy (ganpat):

Volume for cone = 1/3 * pi * r square * h = 1/3 * 3.14 * 3.5 square * 18 = 230. 79 inch cube.. Happy Halloween !!!!! :)